You propose a new temperature scale with temperatures given in °E. You define 0.0°E to
be the normal melting point of ethanol and 255°E to be the normal boiling point of ethanol.
(a) What is the normal boiling point of water in °E? (b) A temperature change of 10.0E°
corresponds to how many C°? (final Answer: a) 284.2°E, b) 7.5C°

please show me the complete working solution and rationale.
Answer #1

M.P of ethanol = -114 C and BP of ethanol = 78.37 C

according to new scale

-114 Celsius corresponds to 0 E and 78.37 Celsius corresponds to 255 E

We can use the equation y = mx +C

0 = m * (-114) +C --- (1)

255 = m (78.37) + C ---- (2)

on solving the two equations

we get m = 1.325 and C = 151.11

a) Normal Boiling point of water = 1.333 * 100 + 151.11 = 133.3 + 151.11 = 284.4 E

and 10 E corresponds to 10 = 1.33 * ( T2-T1)

T2 - T1 = 7.5 C